Consider the structure below for the [EOF2 ]2- ion. Note: E represents an unidentified element.

(a) What is the formal charge on E?     ____________

(b) What is the oxidation state of E?   ____________

(c) Which element could E represent?
Answer #1

Ans :-

(a). Formal charge (F.C) on an element = Number of valence electrons (V) - Number of shared electrons (S)/2 - Number of unshared electrons.

So,

F.C on E = 8 - 8/2 - 4

= 8 - 4 - 4

= 8 - 8

= 0

Hence, formal charge on E = 0

(b).

Let oxidation number of E = y

Oxidation number of E + 1 x Oxidation number of O + 2 x Oxidation number of F = - 2

y + 1(-2) + 2(-1) = -2

y - 2 - 2 = -2

y - 4 = - 2

y = - 2 + 4

y = +2

Hence, Oxidation number or state of E = +2

(c). E element could be Xenon (Xe), which contains 8 electrons in its valence shell.